mal 1 No Sa Heading 1 Heading 2 Tae Subtitle Subtle Em... Emphasis 3 incomestatement sis to complete the folewing questions im and Beth Buter are abort to consruct ther income and expense statement fer ending December 31, 2014. They have pur together the following infcrmation for as of December 31, 2014 Beths Salary rear Fims Salary Interest on $47,000 $18,000 savings account Bonds of Beta Corporation 110 70 5,804 1,256 135,302 9,600 285 4,056 458 863 728 650 280 2,700 250 2,200 200 14,000 4,980 10,905 2,530 4,936 9,100 3,250 4,973 Food Personal care Primary residence mortgage Home mortgage payment Homeowners insurance Property Tax Electricity&Water Telephone Cable & Gas and auto maintenance internet magazines Clothing and other miscellianeous expense Medical expenses ost of furniture purchased with charge card on ian 1, 2014 Amount paid on 012 Toyota Corolla Auto loan payment student loan balance-Beth Student loan payments-Beth New student loan to cover tuitions and books-im Federal income taxes tate income taxes Social security and Medicare Health and disability insurance Gift and donation Family vacation Auto nsurance Contribution to emergency tund sevings contribution to retirement savings contribution to the checking account e by Dec 1, 2034 2,592 3,760 3,200 850 1200 1600 15 using the nformation provided hint some information are useless, prepare an income and expense statement for the Sutiers for the Butlers financ al conditions regarding (a) savings ratic and ibi debt service ratio year encing December 31, 2014 Comment on the
Answer #1
Income and Expense Statement $
    1 Income
Salaries (Jim+ Beth)                 65,000
Interests on Savings account and Bonds                       180
Total Income                 65,180
    2 Expense
Food                   5,804
Personal Care                   1,256
Home mortgage Payment                   9,600
House Insurance                       285
Property tax                   4,056
Electricity & Water                       458
Telephone                       863
Cable & Internet                       728
Gas and Auto maintenance                       650
Books and Magazine                       280
Clothing and other miscellaneous                   2,700
Medical                         250
Auto loan payment                   4,980
Furniture payment                       200
Student Loan payment                   2,530
Federal Taxes                   9,100
State Income taxes                   3,250
Social Security and Medicare                   4,973
Health and disability insurance                   2,592
Auto Insurance                       850
Total Expenses                 55,405
Surplus                   9,775

Savings of Jim and Beth are $ 9775 -3760 = $ 6015.

a ) Savings Ratio = Savings / Total Income = 6015 / 65180 = 9.23%.

They are able to save 9.23% of their annual income, which is a very healthy sign of their financial condition.

b) Debt service Ratio = Payments made pertaining to debt / earnings.

payments pertaining to student loan + mortgage + furniture+ auto loan = 4980+200+2530+9600 = 17,130\-

Debt service Ratio = 17130 / 65180 = 26.6%